# Generating Fine Details of Entity Interactions

### v1

@@ B.3.2 Prompt Completion
Come up with a description of two animals doing [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {”topic”: [content], ”prompt”: description} The description must contains the exact [content] word. For example: Concept: ”High-Fiving” Description: A dolphin and a seal leap from the water, high-fiving with their flippers. Then, the output should be: {”topic”: ”High-Fiving”, ”prompt”: ”A dolphin and a seal leap from the water, high-fiving with their flippers.”}
@@ Appendix C Model Evaluation Details
In this section, we first present automatic evaluation by scenario based on the VLM rating, ImageReward [29], CLIPScore [7] and BLIP-VQA [12] (Table 5), and then provide running examples for comparison of human evaluation and automatic evaluation.
Functional Relation Compositional Relation Multi-subject Interaction
GPT-4o ImReward CLIPS. B-VQA GPT-4o ImReward CLIPS. B-VQA GPT-4o ImReward CLIPS. B-VQA
SD 4.183 1.471 0.914 0.430 4.245 0.949 0.851 0.184 3.740 1.247 0.917 0.203
+ GPT Rewrite 4.027 1.285 0.881 0.323 4.250 0.902 0.846 0.171 3.775 1.206 0.910 0.198
+ GPT Refine 4.085 1.401 0.889 0.376 4.035 0.852 0.830 0.169 3.705 1.216 0.904 0.204
+ Multi-seed 4.220 1.502 0.935 0.472 4.250 1.019 0.872 0.197 3.720 1.242 0.934 0.210
DALL·E 3 4.500 1.343 0.869 0.392 4.690 0.938 0.824 0.190 4.290 1.139 0.868 0.193
DetailScribe 4.650 1.598 0.936 0.482 4.580 1.123 0.875 0.216 4.255 1.378 0.935 0.242
@@
C.1 Agreement between human evaluation and automatic Evaluation
We have used the GPT-4o model as a VLM-based automatic evaluation metric for comparing different models. The GPT-4o model takes the same instruction as our human evaluators and directly outputs a score for the individual images. We further adopted ImageReward  [29], CLIPScore [7] and BLIP-VQA [12] as our automatic evaluator. In this section, we present detailed examples of automatic judgments on generated images from different models to illustrate their capabilities and limitations in evaluating complex concepts involving multi-entity interactions. Specifically, in Fig. 7 we include two example where the automatic evaluator failed to recognize the incorrect intersections between objects or missing key components in the prompt, leading to uniformly high scores for all generated images. Additionally, we provide another example where most evaluators gave consistent ratings.
Capable to align with human in most evaluations, VLM-based evaluator tends to give high scores. These examples highlight the current challenges in using GPT-4o as a judge for complex compositional reasoning tasks. We recommend human evaluation for future experiment if cost and throughput allows.
Figure 7: Example of automatic evaluation misinterpreted and successfully interpreted the critical interaction in prompts. The details in the prompt missed by the evaluators are highlighted in blue with an underline. (H: human, G: GPT-4o, I: ImageReward, C: CLIPScore, B: BLIP-VQA, Highest score of each metrics are highlighted in red)
C.2 VLM Evaluation Prompt Implementation Details
The above images were generated with the prompt: [prompt]. Please rate text-image alignment score of each image from 1 to 5, focus on [topic] and follow the criteria: 1: poor interaction, subject(s) not acting correctly, 2: subject(s) incorrect/inaccurate 3: critical part missing (e.g. missing critical tools or patterns to complete [topic]), 4: nearly perfect but some subparts need further improvement (e.g. needs to refine appearance of tools or limbs), 5: image perfectly depicts [prompt]. Return the score in angle brackets <<<>>>. For example, if the image is nearly perfect and got score 4, response: $<$4$>$
C.2.2 Compositional Spatial Relationships
You are my assistant to identify objects and their spatial layout in the image. According to the image, evaluate if the [prompt] is correctly portrayed in the image. Give a score from 1 to 5 according the criteria: 5: correct spatial layout ([topic]) in the image for all objects mentioned in the text. 4: basically, spatial layout of objects matches the text. 3: spatial layout not aligned properly with the text. 2: image not aligned properly with the text. 1: image almost irrelevant to the text. Return the score in angle brackets <<<>>>. For example, if the image’s spatial layout of objects matches the text and got score 4, response: $<$4$>$
@@ Appendix D Additional Results and Analysis
D.1 Qualitative Examples
Fig. 8 provides additional qualitative examples generated by DetailScribe and other baselines. Overall, DetailScribe is capable of generating high-fidelity, realistic, and faithful images according to different complex language descriptions.
@@ D.2 Qualitative Studies on the Effect of Concept Decomposition and Progressive Refinement
We present a comparative analysis of VLM-generated critiques and refinements with and without the explicit concept decomposition step. Fig. 9 illustrates the result. The results indicate that incorporating concept decomposition significantly enhances the VLM’s ability to focus on meaningful entity interactions rather than overemphasizing fine-grained image details, such as minor variations in subject expressions.
@@ D.3 Qualitative Studies on the Effect of Global Seeds
A current limitation of the DetailScribe framework lies in its reliance on the assumption that the initial image generated without prompt refinement has a roughly correct global scene structure. This limitation arises from the inability of the re-denoising step to introduce large-scale changes to the image, such as adding missing subjects or significantly altering the scene layout. Our examples in Fig. 10 illustrate that, due to the stochasticity inherent in text-to-image generative models, different global random seeds at inference yield different initial images, which in turn affect the final output of our system. While our approach effectively resolves issues like entity interactions in the image, it struggles with global-scale edits, such as adjusting the global layout or zoom level of the scene. As a potential direction for future work, one can consider sampling multiple initial images simultaneously and performing a post-hoc selection process to identify the most suitable candidates for refinement, leveraging VLM feedback to improve global scene accuracy.
@@ E.1.2 Selected prompts
E.2 Compositional Spatial Relationships
@@ E.2.2 Selected prompts
E.3 Multi-subject Interactions
E.3.1 Statistic of topic
E.3.2 Selected prompts
{adjustwidth}
2em0em
  1. 1.
    Topic: Bouncing
    Prompt:
    A frog and a grasshopper take turns bouncing across lily pads on a pond.
  2. 2.
    Topic: Huddling-for-Warmth
    Prompt:
  3. 3.
    Topic: Jumping-Rope
    Prompt:
  4. 4.
    Topic: High-Fiving
    Prompt:
  5. 5.
    Topic: Dancing
    Prompt:
  6. 6.
    Topic: Lifting-Together
    Prompt:
  7. 7.
    Topic: Balancing-a-Ball
    Prompt:
  8. 8.
    Topic: Digging-Together
    Prompt:
  9. 9.
    Topic: Building-a-Nest
    Prompt:
  10. 10.
    Topic: Sharing-Food
    Prompt:

### v2

@@ B.3.2 Prompt Completion
Come up with a description of a scene which is a novel combination of [content], the description should be similar as the following example and uncommon to be observed. Do not use passive voice. Double check the description to focus on major relation, which is [content]. Write down your answer in this format: {"topic": [content], "prompt": description} For example: Concept: "tic-tac-toe" Description: A tic-tac-toe composed by tomato and cucumber as the players symbols. Then, the output should be: {"topic": "tic-tac-toe", "prompt": "A tic-tac-toe composed by tomato and cucumber as the players symbols."}
@@ Appendix C Model Evaluation Details
In this section, we include the evaluation guidelines for human and MLLM annotators, and then present automatic evaluation by scenario based on the MLLM rating, ImageReward Xu et al. (2023), CLIPScore Hessel et al. (2022) and BLIP-VQA Huang et al. (2023a) (Table 5). We also provide running examples for comparison of human evaluation and automatic evaluation.
C.1 Human evaluation
In the human evaluation process, we asked annotators to rate images on a Likert scale. For each prompt, annotators were presented with images generated by all models, which were randomly shuffled to mitigate order bias. Annotators were instructed to assign a score from 1 to 5 based on image-text alignment, following these guidelines:
  • 1: The image is completely irrelevant to the prompt.
  • 3: The image includes most relevant objects, but some elements implied by the prompt are missing (e.g., missing critical tools or patterns to complete the interaction).
  • 4: The image captures most relevant objects and infers additional ones successfully, but there are minor issues with object relationships (e.g. improvement to appearance of tools, subpart, or limbs needed.)
C.2 MLLM Evaluation Prompt Implementation Details
We prompt the MLLM to assign a score from 1 to 5 based on the following guidelines.
The above images were generated with the prompt: [prompt]. Please rate the text-image alignment score of each image from 1 to 5, focusing on [topic] and follow the criteria: 1: poor interaction, subject(s) not acting correctly. 2: subject(s) incorrect/inaccurate. 3: critical part missing (e.g., missing critical tools or patterns to complete [topic]). 4: nearly perfect but some subparts need further improvement (e.g., needs to refine appearance of tools or limbs, subject is not [topic] correctly). 5: image perfectly depicts [topic]. Return the score in angle brackets <>. For example, if the image is nearly perfect and got score 4, response: <4>
C.2.2 Compositional Spatial Relationships
You are my assistant to identify objects and their spatial layout in the image. According to the image, evaluate if the text [prompt] is correctly portrayed in the image. Give a score from 1 to 5 according to the criteria: 5: correct spatial layout ([topic]) in the image for all objects mentioned in the text. 4: basically, spatial layout of objects matches the text. 3: spatial layout not aligned properly with the text. 2: image not aligned properly with the text. 1: image almost irrelevant to the text. Return the score in angle brackets <>. For example, if the image’s spatial layout of objects matches the text and got score 4, response: <4>
C.2.3 Automatic evaluation Results
Functional Relation Multi-subject Interaction Compositional Relation
Qwen2.5 ImReward CLIPS. B-VQA Qwen2.5 ImReward CLIPS. B-VQA Qwen2.5 ImReward CLIPS. B-VQA
SD 4.271 1.471 0.914 0.430 3.780 1.247 0.917 0.203 4.550 0.949 0.851 0.184
+ GPT Rewrite 4.133 1.285 0.881 0.323 3.735 1.206 0.910 0.198 4.445 0.902 0.846 0.171
+ GPT Refine 4.198 1.401 0.889 0.376 3.755 1.216 0.904 0.204 4.355 0.852 0.830 0.169
+ Multi-seed 4.282 1.502 0.935 0.472 3.845 1.242 0.934 0.210 4.565 1.019 0.872 0.197
DALL·E 3 4.477 1.343 0.869 0.392 4.310 1.139 0.868 0.193 4.850 0.938 0.824 0.190
DetailScribe 4.582 1.598 0.936 0.482 4.275 1.378 0.935 0.242 4.830 1.123 0.875 0.216
@@ C.2.3 Automatic evaluation Results
C.3 Agreement between human evaluation and automatic Evaluation
We have used the Qwen2.5-VL-32B model as an MLLM-based automatic evaluation metric for comparing different models. The Qwen2.5 model takes the instruction as shown in Appendix C.2.2 and directly outputs a score for the individual images. We further adopted ImageReward  Xu et al. (2023), CLIPScore Hessel et al. (2022) and BLIP-VQA Huang et al. (2023a) as our automatic evaluator. In this section, we present detailed examples of automatic judgments on generated images from different models to illustrate their capabilities and limitations in evaluating complex concepts involving multi-entity interactions. Specifically, in Fig. 7 we include two example where the automatic evaluator failed to recognize the incorrect intersections between objects or missing key components in the prompt, leading to uniformly high scores for all generated images. Additionally, we provide another example where most evaluators gave consistent ratings.
Capable to align with human in most evaluations, MLLM-based evaluator tends to give high scores. These examples highlight the current challenges in using MLLM as a judge for complex compositional reasoning tasks. We recommend human evaluation for future experiment if cost and throughput allows.
Figure 7: Example of automatic evaluation misinterpreted and successfully interpreted the critical interaction in prompts. The details in the prompt missed by the evaluators are highlighted in blue with an underline. (H: human, Q: Qwen2.5, I: ImageReward, C: CLIPScore, B: BLIP-VQA, Highest score of each metrics are highlighted in red)
@@ Appendix D Additional Results and Analysis
D.1 Qualitative Examples
Fig. 8 provides additional qualitative examples generated by DetailScribe and other baselines. Overall, DetailScribe is capable of generating faithful images according to different complex language descriptions.
@@ D.2 Qualitative Studies on the Effect of Concept Decomposition and Progressive Refinement
We present a comparative analysis of MLLM-generated critiques and refinements with and without the explicit concept decomposition step. Fig. 9 illustrates the result. The results indicate that incorporating concept decomposition significantly enhances the MLLM’s ability to focus on meaningful entity interactions rather than overemphasizing fine-grained image details, such as minor variations in subject expressions.
@@ D.3 Qualitative Studies on the Effect of Global Seeds
A current limitation of the DetailScribe framework lies in its reliance on the assumption that the initial image generated without prompt refinement has a roughly correct global scene structure. This limitation arises from the inability of the re-denoising step to introduce large-scale changes to the image, such as adding missing subjects or significantly altering the scene layout. Our examples in Fig. 10 illustrate that, due to the stochasticity inherent in text-to-image generative models, different global random seeds at inference yield different initial images, which in turn affect the final output of our system. While our approach effectively resolves issues like entity interactions in the image, it struggles with global-scale edits, such as adjusting the global layout or zoom level of the scene. As a potential direction for future work, one can consider sampling multiple initial images simultaneously and performing a post-hoc selection process to identify the most suitable candidates for refinement, leveraging MLLM feedback to improve global scene accuracy.
@@ E.1.2 Selected prompts
E.2 Multi-subject Interactions
@@ E.2.1 Statistic of topic
E.2.2 Selected prompts
  1. 1.
    Topic: Bouncing
    Prompt:
    A frog and a grasshopper take turns bouncing across lily pads on a pond.
  2. 2.
    Topic: Huddling-for-Warmth
    Prompt:
  3. 3.
    Topic: Jumping-Rope
    Prompt:
  4. 4.
    Topic: High-Fiving
    Prompt:
  5. 5.
    Topic: Dancing
    Prompt:
  6. 6.
    Topic: Lifting-Together
    Prompt:
  7. 7.
    Topic: Balancing-a-Ball
    Prompt:
  8. 8.
    Topic: Digging-Together
    Prompt:
  9. 9.
    Topic: Building-a-Nest
    Prompt:
  10. 10.
    Topic: Sharing-Food
    Prompt:
E.3 Compositional Spatial Relationships
E.3.1 Statistic of topic